# Multimodal Laryngoscopic Video Analysis for Assisted Diagnosis of Vocal Fold Paralysis

### v2

@@ 6.5 Vocal Fold Paralysis Detection
Figure 9: Confusion matrix for the left or right VFP classification on the VFP cases of the SYSU-A dataset.
Figure 10: Confusion matrix for the UVFP classification on the SYSU-A test dataset.
Table 8: We present the results of one-tailed t-tests between various system pairs, along with the corresponding t-values, degrees of freedom (df), and p-values, all at the 95% confidence level. These experiments are performed using the SYSU-A dataset, maintaining a consistent train-test split throughout.
Metric Groups Alternative hypothesis $\mu_{0}<\mu$ t-value df p-value
Accuracy Across modules System 1 vs. System 2 3.96 8 0.0020
System 1 vs. Proposed System 5.00 9 0.0003
System 2 vs. Proposed System 2.50 7 0.0210
Across modalities System 3 vs. Proposed System 6.48 9 0.0001
System 4 vs. Proposed System 4.68 9 0.0005
F-score Across modules System 1 vs. System 2 3.15 9 0.0058
System 1 vs. Proposed System 4.19 9 0.0012
System 2 vs. Proposed System 2.31 7 0.0272
Across modalities System 3 vs. Proposed System 4.53 8 0.0009
System 4 vs. Proposed System 4.37 7 0.0016
Modules vs. VFP Detection Performance
Accurate midline prediction is critical for analyzing left and right vocal fold movements, as it serves as a reference for measuring angle deviations and calculating distances between sample points on the vocal folds and the midline. However, in real-world applications, the laryngeal camera’s position is not always fixed, complicating metric computation. Calibrating the midline by fitting a quadratic function has proven effective for improving midline prediction, leading to better classification performance. The associated results are detailed in Table 7. The system’s native setting serves as a foundational benchmark from which incremental enhancements are realized. The introduction of the proposed quadratic fitting method improves the accuracy from 73% to 78% and also enhances the specificity from 70% to 78%, significantly reducing the risk of misdiagnosing healthy patients. Further improvements are achieved by incorporating diffusion refinement into the glottis segmentation process. As indicated in Table 7, sensitivity grows from 77% to 82% and specificity further increases to 81%. The diffusion refinement is applied only to images where the glottis is not detected by the YOLO-v5 detector (discussed in Section 3.3), as this technique specifically targets false alarms (FAs) in glottis detection (described in Section 6.4). By reducing false alarms, the overall performance of downstream VFP detection improves.
The paired t-test comparing System 1 and System 2 yields a p-value of 0.0020 for accuracy and 0.0058 for the F-score, indicating that quadratic fitting contributes to statistically significant improvement at the 5% significance level. The data also suggest that System 2, when compared to the proposed system, exhibits a p-value of 0.0210 for accuracy and 0.0272 for the F1-score, which underscores the positive effect of the diffusion-model based refinement. Additionally, the evaluation of our proposed system against System 1, characterized by basic settings, reveals profound enhancements in both accuracy and F-score, with p-values of 0.0003 and 0.0005 respectively, thereby affirming the combined efficacy of the proposed modules for VFP detection.To corroborate these findings, our analysis includes additional metrics such as t-values and degrees of freedom (df), further substantiating the significance of our advancements.
Multimodality vs. VFP Detection Performance
We also investigated the effect of multimodal inputs on VFP detection. When the AGA movement metrics were removed, classification performance decreased, with the F-score dropping to 74%, and sensitivity and specificity falling by an absolute value of 8.13% and 10.7%, respectively. These results highlight the importance of integrating both audio and AGA movement modalities for robust performance. Additionally, Table 7 demonstrates the significant contribution of the audio modality in VFP detection, as excluding audio leads to a substantial drop in all evaluation metrics, with decreases of over 10%.
The t-tests conducted between the proposed multimodal system and the single-modality system are detailed in Table 8. When comparing System 3 and System 4 to the proposed system, all performance metrics result in p-values lower than 0.05 and large t-values. This outcome indicates that the enhancements provided by the multimodal system are statistically significant at the 5% significance level and substantiates the efficacy of multimodal integration in VFP detection.
Table 9: Unilateral vocal fold paralysis classification through AGA comparison between left and right vocal folder
Systems Accuracy (%) Precision (%) Recall (%) F-score (%)
System 1 75.60 81.18 75.60 75.87
System 2 78.50 78.87 78.50 78.63
Proposed System 82.37 82.56 82.37 82.44
@@ 6.6 Detection of Unilateral Vocal Fold Paralysis
The classification results on the VFP cases from the SYSU-A dataset are summarized in Table 9, with the corresponding confusion matrix depicted in Fig. 10. With the integration of all enhanced modules, our system achieves impressive performance, with all metrics surpassing the 82% threshold, affirming its proficiency in distinguishing left from right VFP. An ablation study further corroborates these findings. The table illustrates that the adoption of quadratic fitting techniques significantly hones classification accuracy, as reflected by the enhancement in performance metrics. The addition of diffusion-based refinement also contributes to superior classification outcomes, excelling beyond the baseline model across all performance metrics. Moreover, the confusion matrix (Fig. 10) clearly shows the benefits of integrating the two enhanced modules. In comparison to the baseline model, which demonstrates a high rate of misclassification from left to right VFP and a lower rate in the opposite direction, the system predominately predicts right VFP over left VFP. However, this bias is substantially mitigated by the integration of both quadratic fitting and diffusion-based refinement techniques, showing a better classification result.
To further illustrate the efficacy of our system, we replace the oracle VFP labels with predictions from the initial multimodal VFP model (outlined in Section 5.2 and results shown in Section 6.5). Subsequently, in the second phase, we differentiate the system VFP cases into left or right VFP by analyzing the variance between LVFDyn and RVFDyn. The confusion matrices depicted in Fig. 10 present the comprehensive three-category classification outcomes. As depicted in Fig. 10, the majority of the samples were correctly identified, with the largest numbers in the confusion matrices forming their diagonals, showing an accuracy of 81.3%.
@@ 7 Conclusion
In this work, we present the Multimodal Laryngoscopic Video Analyzing System (MLVAS), a novel framework that integrates both audio and image processing for enhanced diagnosis of laryngeal disorders. The system employs an audio-based KWS model to capture the full phonation cycle and a vocal fold image detection module to verify vocal folds’ existence. Through HSV color space analysis, MLVAS can extract stroboscopic sequences. As a result, MLVAS enables the automatic generation of video highlights from raw laryngoscopic footage, thus providing a practical diagnostic tool that removes the need for manual pre-editing. Objective metrics for clinical assessment are obtained by a two-stage segmentation process with an initial U-Net model-based segmentation followed by a diffusion model refinement, significantly reducing false positives caused by the initial U-Net segmentation. MLVAS continues with an unsupervised feature extraction approach, applying quadratic fitting to determine the glottal midline and then calculating the angle deviations for left and right vocal folds for all time frames — termed LVFDyn and RVFDyn. Uniting audio and image features, the system adopts a multimodal method for predicting UVFP, showing its effectiveness on a challenging clinical dataset. Furthermore, MLVAS enhances clinical interpretation with visualizations such as GAW and VFDyn, thus equipping clinicians with precise and efficient tools for UVFP diagnosis.
@@ 8 Acknowledgment
This research is funded by the DKU foundation project ”Interdisciplinary Signal Processing Technologies”.

### v3

@@ 7.1 Detection of Vocal Fold Paralysis
Figure 8: Confusion matrix of the proposed system for the left or right VFP classification on the VFP cases of the SYSU-A dataset. System 1: w/o QF and DR. System 2: w/o DR. The proposed system: w/ QF and DR.
Figure 9: Confusion matrix of the proposed system for the overall subject-level UVFP classification on the SYSU-A test dataset.
Table 8: The classification performance (%) of unilateral vocal fold paralysis classification through VFDyn comparison between left and right vocal folds, extracted by different enhancement modules.
Quadratic-
-fitting
Diffusion-
-refinement
Accuracy Precision Recall F-score Sensitivity Specificity
- - 75.60 81.18 75.60 75.87 75.60 75.59
- 78.50 78.87 78.50 78.63 78.50 78.55
82.37 82.56 82.37 82.44 82.37 82.35
@@ 7.2 Detection of Unilateral Vocal Fold Paralysis
The classification results on the VFP cases from the SYSU-A dataset are summarized in Table 8, with the corresponding confusion matrix depicted in Figure 9. With the integration of all enhanced modules, our system achieves impressive performance, with all metrics surpassing the 82% threshold, affirming its proficiency in distinguishing left from right VFP. An ablation study further corroborates these findings. The table illustrates that the adoption of QF techniques significantly hones classification accuracy, as reflected by the enhancement in performance metrics. The addition of diffusion-based refinement also contributes to superior classification outcomes, excelling beyond the baseline model across all performance metrics. Moreover, the confusion matrix (Figure 9) clearly shows the benefits of integrating the two enhanced modules. In comparison to the baseline model, which demonstrates a high rate of misclassification from left to right VFP and a lower rate in the opposite direction, the system predominately predicts right VFP over left VFP. However, this bias is substantially mitigated by the integration of both QF and diffusion-based refinement techniques, showing a better classification result.
To further illustrate the efficacy of our system, we replace the oracle VFP labels with predictions from the initial multimodal VFP model (outlined in Section 5 and results shown in Section 7.1). Subsequently, in the second phase, we differentiate the system VFP cases into left or right VFP by analyzing the variance between LVFDyn and RVFDyn. The confusion matrices depicted in Figure 9 present the comprehensive three-category classification outcomes. The development of the multimodal VFP model, we randomly make train-test splits, yielding 445 cases (81 normal, 364 VFP) for training the detection model, and 75 cases (25 normal, 25 left VFP, 25 right VFP) for testing. As depicted in Figure 9, the majority of the samples were correctly identified, with the largest numbers in the confusion matrices forming their diagonals, showing an accuracy of 81.3%.
@@ 8 Conclusion
In this work, we present the Multimodal Laryngoscopic Video Analyzing System (MLVAS), a novel framework that integrates audio and image processing for the advanced diagnosis of laryngeal disorders. The front-end utilizes an audio-based KWS model to capture the full phonation cycle, along with a vocal fold detection module to confirm the presence of vocal folds. Leveraging HSV color space analysis, MLVAS extracts stroboscopic sequences, enabling the automatic generation of video highlights from raw laryngoscopic footage, thereby eliminating the need for manual pre-editing. The back-end integrates both audio and visual modalities for VFP detection, marking the first application of pre-trained audio models for laryngeal disease diagnosis to encode robust audio features. Visual metrics are computed using a two-stage segmentation pipeline, combining U-Net-based segmentation with diffusion refinement to reduce false positives, followed by quadratic midline fitting to derive LVFDyn and RVFDyn, representing the dynamic angle deviations of the left and right vocal folds. The novel visual metrics enable MLVAS not only to detect VFP but also to distinguish left- from right-sided VFP. By uniting audio and visual features in a cohesive framework, MLVAS achieves superior performance on the clinical dataset and provides interpretable visualizations, such as the GAW and VFDyn, equipping clinicians with reliable tools for precise and efficient UVFP diagnosis.
@@ 9 Acknowledgment
This research is funded by the DKU foundation project ”Interdisciplinary Signal Processing Technologies”. Many thanks for the computational resource provided by the Advanced Computing East China Sub-Center.